The equation of family of curves for which the length of normal is equal to the radius vector

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallvi

let a general point on the curve (x,y)

 length of radius vector = √x2 +y2

 lengt of normal is  =y√(1+(dy/dx)2)

given   √x2 +y2   =y√(1+(dy/dx)2)

   or   x2 +y2   =y2(1+(dy/dx)2)

 or     x2    =y2(dy/dx)2

 or  dy/dx = ±x/y

    ydy  =±xdx

    y2/2  =± x2/2  +c

   y2  ± x2  =c

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin